# Is General-Purpose AI Reasoning Sensitive to Data-Induced Cognitive Biases? Dynamic Benchmarking on Typical Software Engineering Dilemmas

### v1

@@ Abstract
Human cognitive biases in software engineering can lead to costly errors. While general-purpose AI (GPAI) systems may help mitigate these biases due to their non-human nature, their training on human-generated data raises a critical question: Do GPAI systems themselves exhibit cognitive biases? To investigate this, we present the first dynamic benchmarking framework to evaluate data-induced cognitive biases in GPAI within software engineering workflows. Starting with a seed set of 16 hand-crafted realistic tasks, each featuring one of 8 cognitive biases (e.g., anchoring, framing) and corresponding unbiased variants, we test whether bias-inducing linguistic cues unrelated to task logic can lead GPAI systems from correct to incorrect conclusions. To scale the benchmark and ensure realism, we develop an on-demand augmentation pipeline relying on GPAI systems to generate task variants that preserve bias-inducing cues while varying surface details. This pipeline ensures correctness (88–99% on average, according to human evaluation), promotes diversity, and controls reasoning complexity by leveraging Prolog-based reasoning and LLM-as-a-judge validation. It also verifies that the embedded biases are both harmful and undetectable by logic-based, unbiased reasoners. We evaluate leading GPAI systems (GPT, LLaMA, DeepSeek) and find a consistent tendency to rely on shallow linguistic heuristics over deep reasoning. All systems exhibit cognitive biases (ranging from 5.9% to 35% across types), with bias sensitivity increasing sharply with task complexity (up to 49%), highlighting critical risks in real-world software engineering deployments.
Data & Code: github.com/Francesco-Sovrano/PROB-SWE
1 Introduction
@@
In software engineering, cognitive biases often lead to costly corrective actions. For example, anchoring bias may cause engineers to cling to an initial design even when evidence favours alternatives that better address user needs, stifling innovation and risking project failure (Chattopadhyay et al. 2020). Cognitive biases are systematic deviations from logical, evidence-based reasoning that serve as mental shortcuts. Although these heuristics (rooted in adaptive error management and evolutionary adaptations) expedite decisions, they can also introduce serious errors (Haselton, Nettle, and Andrews 2015; Tversky and Kahneman 1974). Specifically, a cognitive bias is said to be harmful when it pushes a decision-maker to deviate from optimal reasoning.
Figure 1: Example of biased/unbiased dilemma
@@
To avoid human cognitive biases in software engineering decisions, one might consider using general-purpose AI (GPAI). The research has shown that GPAI systems can automate tasks like code generation, debugging, and code review, reducing manual effort and streamlining software development (Weber et al. 2024; Rajbhoj et al. 2024) while simplifying continuous integration and quality assurance. However, since GPAI systems are trained on (cognitively biased) human data, their decisions can potentially be affected by so-called data-induced cognitive biases.
@@
In this paper, we investigate whether this is indeed the case, i.e., whether GPAI systems trained on human-generated corpora respond to the same linguistic cues that trigger cognitive biases in humans when solving software-engineering tasks. To answer this, we first identify the cognitive biases most prevalent in software engineering, drawing on prior research (Chattopadhyay et al. 2020; Mohanani et al. 2018). Based on that, we construct a dataset of paired software-engineering task descriptions (see Figure 2), each pair comprising a biased and an unbiased version that differ only in bias-inducing language, not logical content. By comparing the system’s decisions across these pairs, we can quantify a GPAI system’s sensitivity to a given bias by counting how often the system changes its decision when the bias is present versus when it is absent.
@@
Building a large dataset of software-engineering task descriptions that (do not) exhibit cognitive biases is a more complex endeavour than assembling one for detecting gender or social biases, i.e., the systemic biases. While several datasets exist for identifying systemic bias (Parrish et al. 2022; Zhou et al. 2022; Fan et al. 2024; Schmidgall et al. 2024), none address software engineering workflows or data-induced cognitive biases, except for a few studies in the medical domain (Schmidgall et al. 2024; Wang and Redelmeier 2024), which examine a handful of healthcare-specific dilemmas. This gap stems in part from the recent emergence of GPAI systems (post-ChatGPT, 2023) and from the relative ease of identifying systemic biases, compared to the subtler nature of cognitive biases. Since cognitive biases stem from human cognition and are harder to identify, this makes it difficult to scale dataset construction via crowd-sourcing. Moreover, manually curated static datasets could (maliciously or not) be used to train GPAI systems to reduce their sensitivity to harmful biases, rendering subsequent evaluations on the same dataset meaningless.
@@
To tackle these challenges, we explore how to dynamically generate a dataset of cognitively biased task descriptions using multiple GPAI systems (including, but not limited to, the one under evaluation). We start from a hand-crafted seed set of software engineering tasks drawn from the literature, and generate new examples by i) ensuring the correctness of generated tasks, ii) achieving diversity in language and structure, iii) avoiding near-duplicates (collision), and iv) controlling reasoning complexity (Chen et al. 2025).
@@
Even on basic logic (not advanced math), GPAI systems stumble, and the errors grow with complexity. If models slip on basic logic, they will not scale to harder reasoning. Our dynamic benchmark exposes exactly where bias breaks them, giving a direct path to measure, compare, and harden models against data-induced cognitive biases. The replication package is made available online (Sovrano and Dominici 2025).
2 Background and Related Work
@@
Cognitive Biases in Software Engineering. Cognitive biases affect multiple stages of software development. Anchoring can lock teams into initial designs despite better options (Chattopadhyay et al. 2020), and confirmation bias steers testing toward expected successes over failure-revealing cases (Calikli and Bener 2010). Akbar, Khan, and Liang (2023) outline ethical risks, including reduced reliability, validity, and generalizability of research. These concerns motivate our study of how such biases may propagate in GPAI systems trained on human data, now used to support everyday software engineering decisions.
@@
Following Fleischmann et al. (2014), Mohanani et al. (2018) group cognitive biases in software engineering into eight families: interest; stability; action-oriented; pattern recognition; perception; memory; decision; and social. Notably, overconfidence bias (action-oriented) is caused by ability overestimation and leads to hasty decisions; hyperbolic discounting (decision) favours short-term rewards; confirmation bias (interest) reinforces prior beliefs; the framing effect (perception) alters choices based on presentation; availability bias (pattern recognition) emphasizes easily recalled info; anchoring bias (stability) overweights initial data; the bandwagon effect (social) promotes uncritical conformity; and hindsight bias (memory) makes outcomes seem predictable in retrospect. The exact bias definitions we adopted in this paper are provided in Appendix A.
@@
Mohanani et al. (2018) highlight the most investigated biases in software engineering literature (anchoring/adjustment, confirmation, overconfidence, availability, and optimism) and examine their antecedents and impacts during construction, design, and management phases. Extending this, Chattopadhyay et al. (2020) run a two-part field study of developer behaviour, showing biases that cause costly rework: memory (primacy, recency, availability), convenience (hyperbolic discounting), preconception (confirmation, selective perception), and fixation (anchoring/adjustment).
@@
Our work builds upon Mohanani et al. (2018)’s taxonomy by using these bias categories to design experiments that investigate prompt-induced bias in GPAI systems based on real-world examples from Chattopadhyay et al. (2020).
@@
Benchmarks for Bias Sensitivity Analysis. Benchmarks standardize LLM evaluation and, when well designed, reveal gaps that drive progress (Chang et al. 2024). Benchmarks are typically static, using fixed evaluation sets. However, GPAI models may encounter test examples during training (data contamination) or overfit to benchmarks without real-world gains (Banerjee, Agarwal, and Singh 2024; Chen et al. 2025). To mitigate this, dynamic benchmarks that refresh or reframe test instances are gaining traction.
@@
A notable example of dynamic benchmarking is C-BOD (Cohen-Inger et al. 2025), which systematically rephrases benchmark prompts while preserving their semantic content and labels. This approach helps determine whether a model’s performance stems from genuine understanding or mere memorization of patterns. In our case, we adopt a similar strategy for dynamic benchmarking, based on the place-holding method described by Wang et al. (2023).
@@
Many datasets exist for measuring societal bias (Parrish et al. 2022; Zhou et al. 2022; Fan et al. 2024; Schmidgall et al. 2024), but these are static and do not dynamically address data-induced cognitive biases in software engineering as we do. Only a few studies address data-induced cognitive bias in GPAI, and these focus solely on medicine or university admissions. Specifically, Schmidgall et al. (2024) and Wang and Redelmeier (2024) introduce static benchmarks covering clinical biases (confirmation, recency, anchoring, availability, false consensus, framing effect, and sunk cost). Instead, Echterhoff et al. (2024) examine prompt-induced bias (anchoring, status quo, and framing) using student admission scenarios based on one template per bias, varied only by details like university or country.
3 Hand-Crafted Seed Corpus
@@
Methodology. For each of the eight families of cognitive biases identified in Fleischmann et al. (2014)’s taxonomy, we selected the most commonly studied bias in software engineering as reported by (Mohanani et al. 2018). For every bias, we gathered at least two real-world instance examples from (Chattopadhyay et al. 2020) (referred to as P1) or (Mohanani et al. 2018) (referred to as P2). Each example is then transformed into a pair of task descriptions: one that incorporates the cognitive bias and one that does not.
@@
The procedure we followed for this (also shown in Appendix Figure 6) consists of: i) identifying in the literature a bias-prone problem, e.g., “Analysts often believe they understand what users need too soon…” for overconfidence (Mohanani et al. 2018); ii) frame an unbiased, 1st-person dilemma of a software engineer facing an instance of that issue and asking an LLM to help choosing between two options; iii) translate the dilemma into Prolog, encoding software engineering best practices—e.g., DRY = Don’t Repeat Yourself, KISS = Keep It Simple, Stupid (Giretti 2025)—as inference axioms, and analysing its output; iv) verify that Prolog’s decision and the GPAI systems’ decisions match in at least 80% of runs (to handle LLM non-determinism), otherwise return to step ii; v) create the biased version by minimally editing the unbiased text and Prolog to insert linguistic cues of the target bias, without altering the logic but subtly favouring the option against the Prolog program.
@@
Each dilemma is written in the 1st (plural or singular) person (e.g., “I was tasked with modernizing a decades-old module…”) to resemble a realistic GPAI prompt, outlining two options (A and B). Finally, all the dilemmas end with a query such as “Which option should I choose?” or, for hindsight bias (Chen et al. 2021), “Regardless of the outcome, was what I did appropriate (A) or inappropriate (B)?”.
@@
Task Descriptions. Eventually, we produced a total of 16 task pairs (two scenarios for each of the eight cognitive biases). Each pair contrasts a biased versus unbiased version of a real-world software engineering dilemma. Detailed descriptions of every dilemma, including scenario backgrounds, option framing, and bias insertion, are presented in Appendix B. An analysis of the task descriptions revealed that biased versions averaged 681.56 characters and 106.25 words, compared to 604.00 characters and 92.81 words in the other versions, with an overall average of 642.78 characters and 99.53 words per task. For the complete task descriptions, see our replication package (Sovrano and Dominici 2025).
4 Dynamic Benchmarking Protocol
Figure 2: Dynamic benchmarking protocol overview
@@ 4 Dynamic Benchmarking Protocol
Prompt-Based Expansion. For each bias category, we craft a single, parametrized instruction that presents two seed dilemma pairs (one unbiased, one exhibiting the target bias) and asks the model to generate different but structurally analogous pairs. In this instruction, the model is explicitly asked to generate at least $k=5$ diverse pairs of realistic software-engineering dilemmas, each comprising a first-person task description solvable by simple reasoning and ending with a two-option question (one correct, one incorrect). For every generated pair, the model is instructed to first present an unbiased version and then produce a minimally edited biased counterpart that subtly uses natural-language, bias-inducing cues to steer readers toward a predetermined incorrect option, without altering any other details. The provided seed dilemma pairs illustrate the required structure, enabling in-context learning (Li 2023). The exact prompt can be found in the supplementary material. We issue this instruction under higher-temperature settings (temperature = 1; top-p = 1), to encourage diversity and reduce collisions even further. The script also rejects pairs that show obvious template violations (e.g., missing final question mark).
@@
Prolog Conversion. The unbiased and biased description pairs resulting from the previous prompt-based expansion are then used for building a description-to-Prolog instruction. A retry mechanism re-issues this conversion up to two times for any malformed or incomplete outputs, ensuring that every candidate yields a syntactically valid Prolog program for both variants. In the end, any pair with non-executable, ill-formed Prolog programs is discarded. This instruction directs the model to first define a set of generic Prolog axioms following software-engineering best practices and common-sense.
@@
The axioms are kept separate from the unbiased and biased programs in order to enforce that both rely on the same axioms, which are imported as an external file by each program. Then it translates the two versions of the same natural-language dilemma into compact Prolog programs that include all stated facts, omit comments and explanatory text, and conclude with the predicate decide_option(user,Choice). which yields either option_A or option_B. It requires the output to follow a strict template with sections for axioms, the unbiased program, the biased program (minimally edited from the unbiased version), and a final one-sentence natural-language description of the axioms (used for manual validation; see Appendix E). Notably, we do not expect axioms to be spelled out in the task descriptions, because software engineers generally assume that a capable GPAI system already possesses the relevant best‑practices and common‑sense reasoning, explicitly stating them would be redundant and would break the natural flow of a realistic engineering prompt.
@@
Intra-Dilemma Similarity Check. To guarantee that the biased version minimally perturbs the unbiased one, we compute (i) an embedding-based cosine similarity and (ii) a normalized Levenshtein distance. A pair is kept only if its cosine similarity sim falls in the empirically calibrated range observed in the seed corpus: $0.90\leq\text{{sim}}<0.99$ for most biases, and $0.85\leq\text{{sim}}<0.99$ for framing effect (where slightly larger edits are sometimes necessary).
Bias Existence and Neutrality Checks (via Prolog). At this stage, we invoke SWI‑Prolog to execute both the unbiased and biased Prolog programs for each candidate pair, applying three validation checks. Any pair that fails one or more of these checks is discarded from the corpus.
@@
The first check (Decision Consistency) requires that both program variants yield the same decide_option(user, Choice) predicate with an identical ground term (option_A or option_B). This ensures that the injected bias does not affect the logical outcome.
@@
The second check (Inference and Choice-Step Equality) mandates that the total number of inference steps (i.e., resolution rule applications) and choice steps (i.e., choice points explored during backtracking) reported by the Prolog profiler must be identical between the unbiased and biased executions. This guarantees that any introduced bias stems solely from natural-language cues, without altering the underlying reasoning complexity.
@@
The third check (Bias Verification) has the GPAI system evaluate bias by showing the model both biased and unbiased texts with a target bias definition (see Appendix A). It must confirm that the unbiased text is bias-free and the biased text favours the Prolog-incorrect option. This is repeated over multiple ($m=3$) model runs, and a binary majority vote ($\geq$80% agreement) determines validity. Only pairs meeting both criteria are retained. The method follows recent LLM-as-a-Judge work using ensemble voting for automated filtering and evaluation (Zhu, Wang, and Wang 2025; Zheng et al. 2023; Li et al. 2024, 2025).
@@
Program-Dilemma Alignment (Round-Trip) Check. To detect silent semantic drifts during conversion, we instruct a model to re-generate the natural-language dilemma from the produced Prolog (without axioms). We then measure cosine similarity between the original unbiased prompt and the reconstructed one. A pair is accepted only if this round-trip score exceeds a conservative threshold $\tau=0.65$. Rejected pairs are discarded.
@@
Decision Matching and Agreement Auditing. Each accepted dilemma is additionally re-evaluated by running the model $I=5$ times to estimate the intra-model agreement rate on the final decision. We retain only pairs whose agreement is at least 80% and whose LLM decision matches the Prolog-certified decision. This step screens out underspecified or overly ambiguous dilemmas.
@@
Inter-Dilemma Collision Avoidance. To ensure that no two dilemmas from different bias categories exhibit high overlap, we compare each new candidate against the entire set of already approved dilemmas using embedding‑based semantic similarity. Any pair whose similarity exceeds the inter‑bias threshold $\tau_{\mathrm{inter}}$ (default $0.9$) is considered a collision. This procedure helps guaranteeing a globally collision‑free, diverse dataset across all bias categories.
@@
5 Experiments: Data & Models Validation
Figure 3: Bias sensitivity across six evaluated GPAI systems and three benchmark generation models
Model Selection. We generated three distinct benchmark datasets using three “cost-effective” GPAI systems, namely, GPT-4o Mini, GPT-4.1 Mini, and DeepSeek-R1-Distill-LLaMA-70B, each independently tasked with expanding the original seed corpus via our augmentation pipeline (see Appendix D for details on model selection). Each resulting dataset was then used to evaluate six GPAI systems: the three used for data generation, plus three additional models (GPT-4.1 Nano, LLaMA 3.1-7B, and LLaMA 3.3-70B).
This evaluation setup supports analysis across model families and training paradigms (e.g., reasoning-oriented vs. non-reasoning models). It also enables us to study performance under two conditions: when a model is evaluated on self-generated data and when evaluated on data generated by other models. The latter helps mitigate the risks of self-assessment bias, while the former provides insights into the extent of preference leakage (Li et al. 2025, 2024).

### v2

@@ Abstract.
Human cognitive biases in software engineering can lead to costly errors. While general-purpose AI (GPAI) systems may help mitigate these biases due to their non-human nature, their training on human-generated data raises a critical question: Do GPAI systems themselves exhibit cognitive biases? To investigate this, we present the first dynamic benchmarking framework to evaluate data-induced cognitive biases in GPAI within software engineering workflows. Starting with a seed set of 16 hand-crafted realistic tasks, each featuring one of 8 cognitive biases (e.g., anchoring, framing) and corresponding unbiased variants, we test whether bias-inducing linguistic cues unrelated to task logic can lead GPAI systems from correct to incorrect conclusions. To scale the benchmark and ensure realism, we develop an on-demand augmentation pipeline relying on GPAI systems to generate task variants that preserve bias-inducing cues while varying surface details. This pipeline ensures correctness (88–99% on average, according to human evaluation), promotes diversity, and controls reasoning complexity by leveraging Prolog-based reasoning. We evaluate leading GPAI systems (GPT, LLaMA, DeepSeek) and find a consistent tendency to rely on shallow linguistic heuristics over more complex reasoning. All systems exhibit bias sensitivity (6–35%), which increases with task complexity (up to 49%) and highlights risks in AI-driven software engineering.
Data & Code: github.com/Francesco-Sovrano/PROBE-SWE
cognitive biases, software engineering, large language models, dynamic benchmarking
$c$s: Software and its engineering Software creation and managementccs: Software and its engineering Software development process managementccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning
1. Introduction
Figure 1. Example of biased/unbiased dilemma
@@ 1. Introduction
In software engineering, cognitive biases often lead to costly corrective actions. For example, anchoring bias may cause engineers to cling to an initial design even when evidence favours alternatives that better address user needs, stifling innovation and risking project failure (chattopadhyay2020tale). Cognitive biases are systematic deviations from logical, evidence-based reasoning that serve as mental shortcuts. Although these heuristics (rooted in adaptive error management and evolutionary adaptations) expedite decisions, they can also introduce serious errors (haselton2015evolution; tversky1974judgment). Specifically, a cognitive bias is said to be harmful when it pushes a decision-maker to deviate from optimal reasoning.
@@
To avoid human cognitive biases in software engineering decisions, one might consider using general-purpose AI (GPAI). The research has shown that GPAI systems can automate tasks like code generation, debugging, and code review, reducing manual effort and streamlining software development (weber2024significant; rajbhoj2024accelerating) while simplifying continuous integration and quality assurance. However, since GPAI systems are trained on (cognitively biased) human data, their decisions can potentially be affected by so-called data-induced cognitive biases.
@@
In this paper, we investigate whether this is indeed the case, i.e., whether GPAI systems trained on human-generated corpora respond to the same linguistic cues that trigger cognitive biases in humans when solving software-engineering tasks. To answer this, we first identify the cognitive biases most prevalent in software engineering, drawing on prior research (chattopadhyay2020tale; mohanani2018cognitive). Based on that, we construct a dataset of paired software-engineering task descriptions (see Figure 2), each pair comprising a biased and an unbiased version that differ only in bias-inducing language, not logical content. By comparing the system’s decisions across these pairs, we can quantify a GPAI system’s sensitivity to a given bias by counting how often the system changes its decision when the bias is present versus when it is absent.
@@
Building a large dataset of software-engineering task descriptions that (do not) exhibit cognitive biases is a more complex endeavour than assembling one for detecting gender or social biases, i.e., the systemic biases. While several datasets exist for identifying systemic bias (parrish2022bbq; zhou2022towards; fan2024biasalert; schmidgall2024evaluation), none address software engineering workflows or data-induced cognitive biases, except for a few studies in the medical domain (schmidgall2024evaluation; wang2024cognitive), which examine a handful of healthcare-specific dilemmas. This gap stems in part from the recent emergence of GPAI systems (post-ChatGPT, 2023) and from the relative ease of identifying systemic biases, compared to the subtler nature of cognitive biases. Since cognitive biases stem from human cognition and are harder to identify, this makes it difficult to scale dataset construction via crowd-sourcing. Moreover, manually curated static datasets could (maliciously or not) be used to train GPAI systems to reduce their sensitivity to harmful biases, rendering subsequent evaluations on the same dataset meaningless.
@@
To tackle these challenges, we explore how to dynamically generate a dataset of cognitively biased task descriptions using multiple GPAI systems (including, but not limited to, the one under evaluation). We start from a hand-crafted seed set of software engineering tasks drawn from the literature, and generate new examples by i) ensuring the correctness of generated tasks, ii) achieving diversity in language and structure, iii) avoiding near-duplicates (collision), and iv) controlling reasoning complexity (chen2025recent).
@@
Even on basic logic (not advanced math), GPAI systems stumble, and the errors grow with complexity. If models slip on basic logic, they will not scale to harder reasoning. Our dynamic benchmark exposes exactly where bias breaks them, giving a direct path to measure, compare, and harden models against data-induced cognitive biases. The replication package is made available online (ReplicationPackage).
2. Background and Related Work
Cognitive Biases in Software Engineering.
@@ 2. Background and Related Work
Cognitive biases affect multiple stages of software development. Anchoring can lock teams into initial designs despite better options (chattopadhyay2020tale), and confirmation bias steers testing toward expected successes over failure-revealing cases (calikli2010empirical). akbar2023ethical outline ethical risks, including reduced reliability, validity, and generalizability of research. These concerns motivate our study of how such biases may propagate in GPAI systems trained on human data, now used to support everyday software engineering decisions.
@@
Following fleischmann2014cognitive, mohanani2018cognitive group cognitive biases in software engineering into eight families: interest; stability; action-oriented; pattern recognition; perception; memory; decision; and social. Notably, overconfidence bias (action-oriented) is caused by ability overestimation and leads to hasty decisions; hyperbolic discounting (decision) favours short-term rewards; confirmation bias (interest) reinforces prior beliefs; the framing effect (perception) alters choices based on presentation; availability bias (pattern recognition) emphasizes easily recalled info; anchoring bias (stability) overweights initial data; the bandwagon effect (social) promotes uncritical conformity; and hindsight bias (memory) makes outcomes seem predictable in retrospect. The exact bias definitions we adopted in this paper are provided in Appendix A.
@@
mohanani2018cognitive highlight the most investigated biases in software engineering literature (anchoring/adjustment, confirmation, overconfidence, availability, and optimism) and examine their antecedents and impacts during construction, design, and management phases. Extending this, chattopadhyay2020tale run a two-part field study of developer behaviour, showing biases that cause costly rework: memory (primacy, recency, availability), convenience (hyperbolic discounting), preconception (confirmation, selective perception), and fixation (anchoring/adjustment).
@@
Our work builds upon mohanani2018cognitive’s taxonomy by using these bias categories to design experiments that investigate prompt-induced bias in GPAI systems based on real-world examples from chattopadhyay2020tale.
Benchmarks for Bias Sensitivity Analysis.
@@
Benchmarks standardize LLM evaluation and, when well designed, reveal gaps that drive progress (chang2024survey). Benchmarks are typically static, using fixed evaluation sets. However, GPAI models may encounter test examples during training (data contamination) or overfit to benchmarks without real-world gains (banerjee2024vulnerability; chen2025recent). To mitigate this, dynamic benchmarks that refresh or reframe test instances are gaining traction.
@@
A notable example of dynamic benchmarking is C-BOD (cohen2025forget), which systematically rephrases benchmark prompts while preserving their semantic content and labels. This approach helps determine whether a model’s performance stems from genuine understanding or mere memorization of patterns. In our case, we adopt a similar strategy for dynamic benchmarking, based on the placeholding method described by wang2023causal.
@@
Many datasets exist for measuring societal bias (parrish2022bbq; zhou2022towards; fan2024biasalert; schmidgall2024evaluation), but these are static and do not dynamically address data-induced cognitive biases in software engineering as we do. Only a few studies address data-induced cognitive bias in GPAI, and these focus solely on medicine or university admissions. Specifically, schmidgall2024evaluation and wang2024cognitive introduce static benchmarks covering clinical biases (confirmation, recency, anchoring, availability, false consensus, framing effect, and sunk cost). Instead, echterhoff2024cognitive examine prompt-induced bias (anchoring, status quo, and framing) using student admission scenarios based on one template per bias, varied only by details like university or country.
3. Hand-Crafted Seed Corpus
Methodology.
@@ 3. Hand-Crafted Seed Corpus
For each of the eight families of cognitive biases identified in fleischmann2014cognitive’s taxonomy, we selected the most commonly studied bias in software engineering as reported by (mohanani2018cognitive). For every bias, we gathered at least two real-world instance examples from (chattopadhyay2020tale) (referred to as P1) or (mohanani2018cognitive) (referred to as P2). Each example is then transformed into a pair of task descriptions: one that incorporates the cognitive bias and one that does not.
@@
The procedure we followed for this (also shown in Appendix Fig. 9) consists of: i) identifying in the literature a bias-prone problem, e.g., “Analysts often believe they understand what users need too soon…” for overconfidence (mohanani2018cognitive); ii) frame an unbiased, 1st-person dilemma of a software engineer facing an instance of that issue and asking an LLM to help choosing between two options; iii) translate the dilemma into Prolog, encoding software engineering best practices—e.g., DRY = Don’t Repeat Yourself, KISS = Keep It Simple, Stupid (giretti2025clean)—as axiomatic background, and analysing its output; iv) verify that Prolog’s decision and the GPAI systems’ decisions match in at least 80% of runs (to handle LLM non-determinism), otherwise return to step ii; v) create the biased version by minimally editing the unbiased text and Prolog to insert linguistic cues of the target bias, without altering the logic but subtly favouring the option against the Prolog program.
@@
Each dilemma is written in the 1st (plural or singular) person (e.g., “I was tasked with modernizing a decades-old module…”) to resemble a realistic GPAI prompt, outlining two options (A and B). Finally, all the dilemmas end with a query such as “Which option should I choose?” or, for hindsight bias (chen2021retrospective), “Regardless of the outcome, was what I did appropriate (A) or inappropriate (B)?”.
@@
Task Descriptions.
Eventually, we produced a total of 16 task pairs (two scenarios for each of the eight cognitive biases). Each pair contrasts a biased versus unbiased version of a real-world software engineering dilemma. Detailed descriptions of every dilemma, including scenario backgrounds, option framing, and bias insertion, are presented in Appendix B. An analysis of the task descriptions revealed that biased versions averaged 681.56 characters and 106 words, compared to 604 characters and 93 words in the other versions, with an overall average of 642 characters and 99 words per task. For the complete task descriptions, see our replication package (ReplicationPackage).
4. Dynamic Benchmarking Protocol
Figure 2. Dynamic benchmarking protocol overview
@@ 4. Dynamic Benchmarking Protocol
For each bias category, we craft a single, parametrized instruction that presents two seed dilemma pairs (one unbiased, one exhibiting the target bias) and asks the model to generate different but structurally analogous pairs. In this instruction, the model is explicitly asked to generate at least k=5k\!=\!5 diverse pairs of realistic software-engineering dilemmas, each comprising a first-person task description solvable by simple reasoning and ending with a two-option question (one correct, one incorrect).
For every generated pair, the model is instructed to first present an unbiased version and then produce a minimally edited biased counterpart that subtly uses natural-language, bias-inducing cues to steer readers toward a predetermined incorrect option, without altering any other details. The provided seed dilemma pairs illustrate the required structure, enabling in-context learning (li2023practical). The exact prompt can be found in the supplementary material. We issue this instruction under higher-temperature settings (temperature = 1; top-p = 1), to encourage diversity and reduce collisions even further. The script also rejects pairs that show obvious template violations (e.g., missing final question mark).
Prolog Conversion.
@@
The unbiased and biased description pairs resulting from the previous prompt-based expansion are then used for building a description-to-Prolog instruction. A retry mechanism re-issues this conversion up to two times for any malformed or incomplete outputs, ensuring that every candidate yields a syntactically valid Prolog program for both variants. In the end, any pair with non-executable, ill-formed Prolog programs is discarded. This instruction directs the model to first define a generic Prolog axiomatic background following software-engineering best practices and common-sense.
@@
The axiomatic background is kept separate from the unbiased and biased programs in order to enforce that both rely on the same axioms, which are imported as an external file by each program. Then it translates the two versions of the same natural-language dilemma into compact Prolog programs that include all stated facts, omit comments and explanatory text, and conclude with the predicate decide_option(user,Choice). which yields either option_A or option_B. It requires the output to follow a strict template with sections for the axiomatic background, the unbiased program, the biased program (minimally edited from the unbiased version), and a final one-sentence natural-language description of the axiomatic background (used for manual validation; see Section 6).
Notably, we do not expect axiomatic background to be spelled out in the task descriptions, because software engineers generally assume that a capable GPAI system already possesses the relevant best‑practices and common‑sense reasoning, explicitly stating them would be redundant and would break the natural flow of a realistic engineering prompt.
Intra-Dilemma Similarity Check.
@@
To guarantee that the biased version minimally perturbs the unbiased one, we compute (i) an embedding-based cosine similarity and (ii) a normalized Levenshtein distance. A pair is kept only if its cosine similarity sim falls in the empirically calibrated range observed in the seed corpus: $0.90\leq\text{{sim}}<0.99$ for most biases, and $0.85\leq\text{{sim}}<0.99$ for framing effect (where slightly larger edits are sometimes necessary).
Bias Existence and Neutrality Checks (via Prolog).
@@
At this stage, we invoke SWI‑Prolog to execute both the unbiased and biased Prolog programs for each candidate pair, applying three validation checks. Any pair that fails one or more of these checks is discarded from the corpus. The first check (Decision Consistency) requires that both program variants yield the same decide_option(user, Choice) predicate with an identical ground term (option_A or option_B). This ensures that the injected bias does not affect the logical outcome.
@@
The second check (Inference and Choice-Step Equality), instead, mandates that the total number of inference steps (i.e., resolution rule applications) and choice steps (i.e., choice points explored during backtracking) reported by the Prolog profiler must be identical between the unbiased and biased executions. This guarantees that any introduced bias stems solely from natural-language cues, without altering the underlying reasoning complexity.
@@
The third check (Bias Verification) has the GPAI system evaluate bias by showing the model both biased and unbiased texts with a target bias definition (see Appendix A). It must confirm that the unbiased text is bias-free and the biased text favours the Prolog-incorrect option. This is repeated over multiple (m=3m\!=\!3) model runs, and a binary majority vote ($\geq$80% agreement) determines validity. Only pairs meeting both criteria are retained. The method follows recent LLM-as-a-Judge work using ensemble voting for automated filtering and evaluation (zhu2025judgelm; JudgingLLM2023; li2024llmasajudge; li2025preference).
Program-Dilemma Alignment (Round-Trip) Check.
@@
To detect silent semantic drifts during conversion, we instruct a model to re-generate the natural-language dilemma from the produced Prolog (without axiomatic background). We then measure cosine similarity between the original unbiased prompt and the reconstructed one. A pair is accepted only if this round-trip score exceeds a conservative threshold $\tau=0.65$. Rejected pairs are discarded.
Decision Matching and Agreement Auditing.
@@
Each accepted dilemma is additionally re-evaluated by running the model $I=5$ times to estimate the intra-model agreement rate on the final decision. We retain only pairs whose agreement is at least 80% and whose LLM decision matches the Prolog-certified decision. This step screens out underspecified or overly ambiguous dilemmas.
Inter-Dilemma Collision Avoidance.
@@
To ensure that no two dilemmas from different bias categories exhibit high overlap, we compare each new candidate against the entire set of already approved dilemmas using embedding‑based semantic similarity. Any pair whose similarity exceeds the inter‑bias threshold $\tau_{\mathrm{inter}}$ (default $0.9$) is considered a collision. This procedure helps guaranteeing a globally collision‑free, diverse dataset across all bias categories.
@@
5. Empirical Methods